# Multilevel neural simulation-based inference

### v2

@@ 1 Introduction
For a long time, SBI was dominated by methods such as approximate Bayesian computation (ABC) (Beaumont et al., 2002; Beaumont, 2019), which compared summary statistics of simulations and of the observed data. However, SBI methods using neural networks to approximate likelihoods (Papamakarios et al., 2019; Lueckmann et al., 2019; Boelts et al., 2022), likelihood ratios (Thomas et al., 2022; Durkan et al., 2020; Hermans et al., 2020), or posterior distributions (Papamakarios and Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Radev et al., 2022) are now quickly becoming the preferred approach. These neural SBI methods are often favoured because they allow for amortisation (Zammit-Mangion et al., 2024), meaning that they require a large offline cost to train the neural network, but once the network is trained, the method can rapidly infer parameters for new observations or different priors without requiring additional costly simulations. This is particularly useful when the simulator is computationally expensive, as it reduces the need to repeatedly run simulations for each new inference task, making the overall process significantly less costly.
@@
Nevertheless, the initial training phase of neural SBI methods typically still requires a large number of simulations, preventing their application on computationally expensive (and often more realistic) models which can take hours of compute time for simulating a single data-point. Examples include most tsunami (Behrens and Dias, 2015), wind farm (Kirby et al., 2023), nuclear fusion (Hoppe et al., 2021) and cosmology (Jeffrey et al., 2025) simulators. One avenue to mitigating this issue is multi-fidelity methods (Peherstorfer et al., 2018): we often have access to a sequence of simulators with increasing computational cost and accuracy which we may be able to use to refine existing methods based on a single simulator.
@@
Figure 1: Low- and high-fidelity cosmological simulations from the CAMELS data (Villaescusa-Navarro et al., 2023) studied in Section˜5.4.
@@
This setting is quite common. One example is simulators requiring the numerical solution of ordinary, partial, or stochastic differential equations, where the choice of mesh size or step-size affects both the accuracy of the solution and the computational cost. Another example arises when modelling complex physical, chemical, or biological processes, where low-fidelity simulators can be obtained by neglecting certain aspects of the system. This is exemplified in Figure˜1, where a high-fidelity cosmological simulation includes baryonic astrophysics and thus appear smoother than the low-fidelity simulation.
@@
The key idea behind multi-fidelity methods is to leverage cheaper, less accurate simulations to supplement the more expensive, high-fidelity simulations, ultimately improving efficiency without sacrificing accuracy. This has been popular in the emulation literature since the seminal work of Kennedy and O’Hagan (2000), but applications to SBI are much more recent and include Jasra et al. (2019); Warne et al. (2018); Prescott and Baker (2020, 2021); Warne et al. (2022); Prescott et al. (2024), who proposed multi-fidelity versions of ABC. More recently, (Krouglova et al., 2025) also proposed a multi-fidelity method to enhance neural SBI approximations of the posterior based on transfer learning. However, their method is not supported by any theoretical guarantees, and our numerical experiments in Section˜5.2 will demonstrate that it is sensitive to choices of hyperparameters and Krouglova et al. (2025) unfortunately do not provide any guidance for selecting these.
@@
In this paper, we propose a novel multi-fidelity method which is broadly applicable to neural SBI methods. Taking neural likelihood and neural posterior estimation as the main case-studies, we show that our approach is able to significantly reduce the computational cost of the initial training through the use of multilevel Monte Carlo (Giles, 2015; Jasra et al., 2020) estimates of the training objective. The approach has strong theoretical guarantees; our main result (Theorem˜1) directly links the reduction in computational cost to the accuracy of the low fidelity simulators, and we demonstrate (in Theorem˜2) how to best balance the number of simulations at each fidelity level in the process. Importantly, and unlike Krouglova et al. (2025), our method has no additional hyperparameter which directly relate to fidelity. Our extensive experiments on models from finance, synthetic biology, and cosmology also demonstrate the significant computational advantages provided by our method.
@@ 2.1 Simulation-based inference
Let $\{\mathbb{P}_{\theta}\}_{\theta\in\Theta}$ be a parametric family of distributions on some space $\mathcal{X}\subseteq\mathbb{R}^{d_{\mathcal{X}}}$ parameterised by $\theta\in\Theta\subseteq\mathbb{R}^{d_{\Theta}}$. We assume that this model is available in the form of a computer code, i.e. as a simulator, wherein simulating from $\mathbb{P}_{\theta}$ is straightforward, but the likelihood function $p(\cdot\,|\,\theta)$ associated with $\mathbb{P}_{\theta}$ is intractable. Simulators can be characterised by a pair $(\mathbb{U},G_{\theta})$, where $\mathbb{U}$ is a distribution (typically simple, such as a uniform or a Gaussian) on a space $\mathcal{U}\subseteq\mathbb{R}^{d_{\mathcal{U}}}$ which captures all of the randomness, and $G_{\theta}:\mathcal{U}\mapsto\mathcal{X}$ is a (deterministic) parametric map called the generator. Simulating $x\sim\mathbb{P}_{\theta}$ can be achieved by first simulating $u\sim\mathbb{U}$, and then applying the generator $x=G_{\theta}(u)$. In this paper, we consider Bayesian inference for the parameters $\theta$ of this simulator-based model given independent and identically distributed (iid) data $x_{1:m}^{o}=\{x_{j}^{o}\}_{j=1}^{m}\in\mathcal{X}^{m}$ collected from some data-generating process. Specifically, we are interested in approximating the posterior with density $\pi(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}p(x_{j}^{o}\,|\,\theta)\pi(\theta)$, where $\pi(\theta)$ is the prior density. As introduced below, this can be achieved via a neural SBI method which approximates the likelihood or posterior.
@@
NLEs (Papamakarios et al., 2019; Lueckmann et al., 2019; Boelts et al., 2022; Radev et al., 2023a) are extensions of the synthetic likelihood approach (Wood, 2010; Price et al., 2018) that use flexible conditional density estimators, typically normalising flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021), as surrogate models for the likelihood function associated with $\mathbb{P}_{\theta}$. The surrogate conditional density $q_{\phi}^{\text{NLE}}:\mathcal{X}\times\Theta\rightarrow[0,\infty)$ where $q_{\phi}^{\text{NLE}}(\cdot\,|\,\theta)$ is a density function for each $\theta\in\Theta$ and $\phi\in\Phi\subseteq\mathbb{R}^{d_{\Phi}}$ denotes its learnable parameters, is trained by minimising the negative log-likelihood with respect to $\phi$ on simulated samples. More precisely, let $\{(\theta_{i},x_{i})\}_{i=1}^{n}$ be the training data such that $\theta_{i}\sim\pi$ are realisations from the prior and $x_{i}\sim\mathbb{P}_{\theta_{i}}$ are realisations from the simulator, then $\hat{\phi}_{\text{MC}}:=\operatorname*{arg\,min}_{\phi\in\Phi}\ell_{\text{MC}}^{\text{NLE}}(\phi)$, where $\ell_{\text{MC}}^{\text{NLE}}$ is an empirical (Monte Carlo) estimate of negative expected log-density:
@@
Once the surrogate likelihood is trained, Markov chain Monte Carlo (MCMC) or variational inference methods are used to sample from the (approximate) posterior distribution $\pi_{\text{NLE}}(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}q^{\text{NLE}}_{\hat{\phi}_{\text{MC}}}(x^{o}_{j}\,|\,\theta)\pi(\theta)$. NLEs can therefore be regarded as being partially amortised—the surrogate $q_{\phi}^{\text{NLE}}$ need not be trained for a new observed dataset, however, MCMC needs to be carried out again to obtain the new posterior.
@@
Instead of learning a surrogate likelihood, NPEs learn a mapping $x\mapsto p(\theta\,|\,x)$ from the data to the posterior using conditional density estimators. These are often based on Gaussian mixture network (Bishop, 1994; Papamakarios and Murray, 2016) or normalising flows (Dinh et al., 2014; Papamakarios et al., 2017; Radev et al., 2022). Similar to NLE, the conditional density $q_{\phi}^{\text{NPE}}:\Theta\times\mathcal{X}^{m}\rightarrow[0,\infty)$ is trained by minimising the negative log likelihood with respect to $\phi$ using data $\{(\theta_{i},x_{1:m,i})\}_{i=1}^{n}$ generated by first sampling from the prior $\theta_{i}\sim\pi$ and then the simulator $x_{1:m,i}=(x_{1,i},\ldots,x_{m,i})\sim\mathbb{P}_{\theta_{i}}$:
@@ 2.2 Multilevel Monte Carlo method
This issue can be mitigated by multilevel Monte Carlo (MLMC), which was first proposed by Heinrich (2001); Giles (2008), and more recently reviewed in Giles (2015); Jasra et al. (2020). Suppose we have a sequence of square-integrable functions $f_{l}:\mathcal{Z}\rightarrow\mathbb{R}$ for $l\in\{0,1,\ldots,L\}$ which are approximations of $f$ and which are ordered such that $f_{L}=f$, and both the cost of evaluation $C_{l}$ and the accuracy (or fidelity) of $f_{l}$ increases with $l$. In that case, MLMC consists of expressing $\mathbb{E}_{z\sim\mu}[f(z)]$ through a telescoping sum and approximating each term through MC based on samples $z^{l}_{i}\sim\mu$ for $i=1,\ldots,n_{l}$ and $l=\{0,1,\ldots,L\}$:
@@
$\approx\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}f_{0}(z^{0}_{i})+\sum_{l=1}^{L}\left(\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}\left(f_{l}\left(z^{l}_{i}\right)-f_{l-1}\left(z^{l}_{i}\right)\right)\right)$
@@ 3 Methodology
such that $f_{\phi}^{L}(u_{1:m},\theta):=f_{\phi}(u_{1:m},\theta)$. This sequence of functions gives evaluations of the log conditional density at evaluations of the (approximate) simulator. Recall that the larger the value of $l$, the more accurate (and computationally expensive) such simulations will tend to be. At this point, we can re-express the objective using a telescoping sum as
@@
$\left\{\theta_{i}^{l},u_{1:m,i}^{l},G^{l}_{\theta^{l}_{i}}\big{(}u_{1,i}^{l}\big{)},\dots,G^{l-1}_{\theta^{l}_{i}}\big{(}u_{m,i}^{l}\big{)}\right\}\quad\text{where}\quad\theta_{i}^{l}\sim\pi,u_{1:m,i}^{l}\sim\mathbb{U},$
@@
This corresponds to an MLMC estimator of our original objective in ˜1. The first term, $h_{0}(\phi)$, approximates $\ell(\phi)$, but is biased since it uses $f^{0}_{\phi}$ (i.e. the lowest fidelity simulator) rather than $f^{L}_{\phi}$ (the highest fidelity simulator). The additional terms, $h_{1}(\phi),\ldots,h_{L}(\phi)$, correct this bias by estimating the expected difference between the objectives at consecutive fidelity levels.
Within each term $h_{l}(\phi)$, the functions $f_{\phi}^{l}$ and $f_{\phi}^{l-1}$ are evaluated on the same samples $u_{1:m}^{l}$ and $\theta^{l}$, i.e., they are seed-matched. This ensures that the terms are coupled and highly correlated, which leads to a reduction in variance (Owen, 2013, Ch. 8) since
@@
# Projecting the gradients (only when conflicting)
@@ 4 Theory
We say that $\mu$ is log-concave if it has a density of the form $\exp(-\psi(z))$ for some convex function $\psi:\mathcal{Z}\rightarrow\mathbb{R}$. We recall that for $r\in[1,\infty]$, $d,d^{\prime}\in\mathbb{N}$ and a non-empty, open, connected set $\mathcal{Z}\subseteq\mathbb{R}^{d}$, the space of vector-valued $r$-integrable functions with respect to a probability distribution $\mu$ is given by $L^{r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{L^{r}(\mu)}:=(\int_{\mathcal{Z}}\|g(x)\|_{2}^{r}\mu(dx))^{\nicefrac{{1}}{{r}}}<\infty\}$. For $\tau\in\mathbb{N}$, the corresponding Sobolev space of vector-valued functions of smoothness $\tau$ is given by $W^{\tau,r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{W^{\tau,r}(\mu)}=(\sum_{|\alpha|\leq\tau}\|D^{\alpha}g\|_{L^{r}(\mu)}^{r})^{\nicefrac{{1}}{{r}}}<\infty\}$, where for a multi-index $\alpha\in\mathbb{N}^{d}$, $D^{\alpha}$ is the weak derivative operator corresponding to $\alpha$. Finally, we recall that a function $g$ is locally $K_{\text{Lip}}$-smooth if its gradient is locally Lipschitz continuous with Lipschitz constant $K_{\text{Lip}}>0$; i.e. for all $z,z^{\prime}$ in some open set of $\mathcal{Z}$, we have that $\|\nabla g(z)-\nabla g(z^{\prime})\|_{2}\leq K_{\text{Lip}}\|z-z^{\prime}\|_{2}$.
@@
    The generators satisfy $\left\|G^{l}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}\leq S_{l}\;$ for $l\in\{0,1,\ldots,L\}$.
@@
    One of the following conditions holds:
    • NLE: $x,\theta\mapsto\log q^{\text{\text{NLE}}}_{\phi}(x|\theta)$ is continuously differentiable, locally $K_{\text{Lip}}(\phi)$-smooth and satisfies the growth condition $\|\nabla_{x}\log q^{\text{\text{NLE}}}_{\phi}(x|\theta)\|_{2}\leq K_{\text{grow}}(\phi)(\|(x,\theta)\|_{2}+1)$ for some $K_{\text{Lip}}(\phi),K_{\text{grow}}(\phi)>0$.
    • NPE: $x_{1},\ldots,x_{m},\theta\mapsto\log q^{\text{\text{NPE}}}_{\phi}(\theta|x_{1},\ldots,x_{m})$ is continously differentiable, locally $K_{\text{Lip}}(\phi)$-smooth and satisfies the growth condition $\|\nabla_{\theta}\log q^{\text{\text{NLE}}}_{\phi}(\theta|x_{1},\ldots,x_{m})\|_{2}\leq K_{\text{grow}}(\phi)(\|(x_{1},\ldots,x_{m},\theta)\|_{2}+1)$ for some $K_{\text{Lip}}(\phi),K_{\text{grow}}(\phi)>0$.
@@
$\text{Var}\left[h_{0}(\phi)\right]$ $\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1\right),$
$\text{and}\;\text{Var}\big{[}h_{l}(\phi)\big{]}$ $\leq\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1\right).$
See Section˜A.1 for the proof, and we emphasise that the variance is over both parameters $\theta$ and noise $u_{1:m}$. (A1) requires log-concavity, which is a strong condition, but this is only required of the prior and the base measure. However, in SBI these tend to be simple distributions such as Gaussians or uniforms, which satisfy this assumption; see Saumard and Wellner (2014) for how to verify log-concavity in practice. (A2) is relatively mild: it asks that $G^{0},\ldots,G^{L}$ have at least one derivative in $\theta$ and $u$, and for these generators and their derivatives to have a fourth moment. This will hold when the simulators are used to define sufficiently well-behaved data-generating processes, but will be violated for sufficiently heavy-tailed distributions (e.g. the Cauchy). Finally, (A3) is mild; it holds when the gradient of the log conditional density estimator is Lipschitz continuous, which is for example the case when the conditional density is twice continuously differentiable and the Hessian is bounded (see e.g. Lemma 2.3 of Wright and Recht (2022)). It also holds for models such as mixtures of Gaussians, which are used in mixture density networks and for normalising flows with sufficiently regular transformations; see Table 1 in Liang et al. (2022) for some known Lipschitz continuous transformations. There are two key implications of Theorem˜1. The first is a bound on the variance of the MC objective:
$\text{Var}\left[{\ell}_{\text{MC}}(\phi)\right]\leq\frac{K(\phi)}{n}\left(\left\|G\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1\right),$ (4)
@@
$\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1\right).$ (5)
In order to make this bound small, we need to make each term small. Typically, $\|G^{0}\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}$ will be large, but this will be counter-balanced by taking $n_{0}$ to be large. For the higher-fidelity levels, the number of samples $n_{l}$ will typically be smaller, but this will be counter-balanced by the fact that $\|G^{l}-G^{l-1}\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}$ is small whenever $G^{l-1}$ is a good approximation of $G^{l}$. As we now show, we can directly use ˜5 to get an indication of how to select $n_{0},\ldots,n_{L}$.

### v4

@@ 1 Introduction
For a long time, SBI was dominated by methods such as approximate Bayesian computation (ABC) (Beaumont et al., 2002; Beaumont, 2019), which compared summary statistics of simulations and of the observed data. However, SBI methods using neural networks to approximate likelihoods (Papamakarios et al., 2019; Lueckmann et al., 2019; Boelts et al., 2022), likelihood ratios (Thomas et al., 2022; Durkan et al., 2020; Hermans et al., 2020), or posterior distributions (Papamakarios and Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Radev et al., 2022) are now quickly becoming the preferred approach. These neural SBI methods are often favoured because they allow for amortisation (Zammit‐Mangion et al., 2024), meaning that they require a large offline cost to train the neural network, but once the network is trained, the method can rapidly infer parameters for new observations or different priors without requiring additional costly simulations. This is particularly useful when the simulator is computationally expensive, as it reduces the need to repeatedly run simulations for each new inference task, making the overall process significantly less costly.
@@
Nevertheless, the initial training phase of neural SBI methods typically still requires a large number of simulations, preventing their application on computationally expensive (and often more realistic) models which can take hours of compute time for simulating a single data-point. Examples include most tsunami (Behrens and Dias, 2015), wind farm (Kirby et al., 2023), nuclear fusion (Hoppe et al., 2021) and cosmology (Jeffrey and others, 2025) simulators. One avenue to mitigate this issue is multi-fidelity methods (Peherstorfer et al., 2018): we often have access to a sequence of simulators with increasing computational cost and accuracy which we may be able to use to refine existing methods based on a single simulator.
@@
Figure 1: Low- and high-fidelity cosmological simulations from the CAMELS data (Villaescusa-Navarro and others, 2023) studied in Section˜5.4.
@@
This setting is quite common. One example is simulators requiring the numerical solution of ordinary, partial, or stochastic differential equations, where the choice of mesh size or step-size affects both the accuracy of the solution and the computational cost. Another example arises when modelling complex physical, chemical, or biological processes, where low-fidelity simulators can be obtained by neglecting certain aspects of the system. This is exemplified in Figure˜1, where a high-fidelity cosmological simulation includes baryonic astrophysics and thus appears smoother than the low-fidelity simulation.
@@
The key idea behind multi-fidelity methods is to leverage cheaper, less accurate simulations to supplement the more expensive, high-fidelity simulations, ultimately improving efficiency without sacrificing accuracy. This has been popular in the emulation literature since the seminal work of Kennedy and O’Hagan (2000), but applications to SBI are much more recent and include Jasra et al. (2019); Warne et al. (2018); Prescott and Baker (2020, 2021); Warne et al. (2022); Prescott et al. (2024), who proposed multi-fidelity versions of ABC. More recently, (Krouglova et al., 2025) also proposed a multi-fidelity method to enhance neural SBI approximations of the posterior based on transfer learning. However, their method is not supported by theoretical guarantees.
@@
In this paper, we propose a novel multi-fidelity method which is broadly applicable to neural SBI methods. Taking neural likelihood and neural posterior estimation as the main case studies, we show that our approach is able to significantly reduce the computational cost of the initial training through the use of multilevel Monte Carlo (Giles, 2015; Jasra et al., 2020) estimates of the training objective. The approach has strong theoretical guarantees; our main result (Theorem˜1) directly links the reduction in computational cost to the accuracy of the low-fidelity simulators, and we demonstrate (in Theorem˜2) how to best balance the number of simulations at each fidelity level in the process. Our extensive experiments on models from finance, synthetic biology, and cosmology also demonstrate the significant computational advantages provided by our method.
@@ 2.1 Simulation-based inference
Let $\{\mathbb{P}_{\theta}\}_{\theta\in\Theta}$ be a parametric family of distributions on some space $\mathcal{X}\subseteq\mathbb{R}^{d_{\mathcal{X}}}$ parameterised by $\theta\in\Theta\subseteq\mathbb{R}^{d_{\Theta}}$. We assume that this model is available in the form of a computer code, i.e. as a simulator, where simulating from $\mathbb{P}_{\theta}$ is straightforward, but the likelihood function $p(\cdot\,|\,\theta)$ associated with $\mathbb{P}_{\theta}$ is intractable. Simulators can be characterised by a pair $(\mathbb{U},G_{\theta})$, where $\mathbb{U}$ is a distribution (typically simple, such as a uniform or a Gaussian) on a space $\mathcal{U}\subseteq\mathbb{R}^{d_{\mathcal{U}}}$ which captures all of the randomness, and $G_{\theta}:\mathcal{U}\mapsto\mathcal{X}$ is a (deterministic) parametric map called the generator. Simulating $x\sim\mathbb{P}_{\theta}$ can be achieved by first simulating $u\sim\mathbb{U}$, and then applying the generator $x=G_{\theta}(u)$. In this paper, we consider Bayesian inference for the parameters $\theta$ of this simulator-based model given independent and identically distributed (iid) data $x_{1:m}^{o}=\{x_{j}^{o}\}_{j=1}^{m}\in\mathcal{X}^{m}$ collected from some data-generating process. Specifically, we are interested in approximating the posterior with density $\pi(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}p(x_{j}^{o}\,|\,\theta)\pi(\theta)$, where $\pi(\theta)$ is the prior density. As introduced below, this can be achieved via a neural SBI method which approximates the likelihood or posterior.
@@
NLEs (Papamakarios et al., 2019; Lueckmann et al., 2019; Boelts et al., 2022; Radev et al., 2023a) are extensions of the synthetic likelihood approach (Wood, 2010; Price et al., 2018) that use flexible conditional density estimators, typically normalising flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021), as surrogate models for the likelihood function associated with $\mathbb{P}_{\theta}$. The surrogate conditional density $q_{\phi}^{\text{NLE}}:\mathcal{X}\times\Theta\rightarrow[0,\infty)$ where $q_{\phi}^{\text{NLE}}(\cdot\,|\,\theta)$ is a density function for each $\theta\in\Theta$ and $\phi\in\Phi\subseteq\mathbb{R}^{d_{\Phi}}$ denotes its learnable parameters, is trained by minimising the negative log-likelihood with respect to $\phi$ on simulated samples. More precisely, let $\{(\theta_{i},x_{i})\}_{i=1}^{n}$ be the training data such that $\theta_{i}\sim\pi$ are realisations from the prior and $x_{i}\sim\mathbb{P}_{\theta_{i}}$ are realisations from the simulator. Then $\hat{\phi}_{\text{MC}}:=\operatorname*{arg\,min}_{\phi\in\Phi}\ell_{\text{MC}}^{\text{NLE}}(\phi)$, where $\ell_{\text{MC}}^{\text{NLE}}$ is an empirical (Monte Carlo) estimate of negative expected log-density:
@@
Once the surrogate likelihood is trained, Markov chain Monte Carlo (MCMC) or variational inference methods are used to sample from the (approximate) posterior distribution $\pi_{\text{NLE}}(\theta\,|\,x_{1:m}^{o})\propto\prod_{j=1}^{m}q^{\text{NLE}}_{\hat{\phi}_{\text{MC}}}(x^{o}_{j}\,|\,\theta)\pi(\theta)$. NLEs can therefore be regarded as being partially amortised—the surrogate likelihood need not be trained for a new observed dataset, however, MCMC needs to be carried out again to obtain the new posterior.
@@
Instead of learning a surrogate likelihood, NPEs learn a mapping $x\mapsto p(\theta\,|\,x)$ from the data to the posterior using conditional density estimators. These are often based on mixture density networks (Bishop, 1994; Papamakarios and Murray, 2016) or normalising flows (Dinh et al., 2014; Papamakarios et al., 2017; Radev et al., 2022). Similar to NLE, the conditional density $q_{\phi}^{\text{NPE}}:\Theta\times\mathcal{X}^{m}\rightarrow[0,\infty)$ is trained by minimising the negative log likelihood with respect to $\phi$ using data $\{(\theta_{i},x_{1:m,i})\}_{i=1}^{n}$ generated by first sampling from the prior $\theta_{i}\sim\pi$ and then the simulator $x_{1:m,i}=(x_{1,i},\ldots,x_{m,i})\sim\mathbb{P}_{\theta_{i}}$:
@@ 2.2 Multilevel Monte Carlo method
This issue can be mitigated by multilevel Monte Carlo (MLMC), which was first proposed by Heinrich (2001); Giles (2008), and more recently reviewed in Giles (2015); Jasra et al. (2020). Suppose we have a sequence of square-integrable functions $f_{l}:\mathcal{Z}\rightarrow\mathbb{R}$ for $l\in\{0,1,\ldots,L\}$ which are approximations of $f$ and which are ordered such that $f_{L}=f$, and both the cost of evaluation $C_{l}$ and the accuracy (or fidelity) of $f_{l}$ increase with $l$. In that case, MLMC consists of expressing $\mathbb{E}_{z\sim\mu}[f(z)]$ through a telescoping sum and approximating each term through MC based on samples $z^{l}_{i}\sim\mu$ for $i=1,\ldots,n_{l}$ and $l\in\{0,1,\ldots,L\}$:
@@
$\approx\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}f_{0}(z^{0}_{i})+\sum_{l=1}^{L}\left(\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}\left(f_{l}\left(z^{l}_{i}\right)-f_{l-1}\left(z^{l}_{i}\right)\right)\right).$
@@ 3 Methodology
such that $G^{L}_{\theta}(u):=G_{\theta}(u)$ and $f_{\phi}^{L}(u_{1:m},\theta):=f_{\phi}(u_{1:m},\theta)$. This sequence of functions gives evaluations of the log conditional density at evaluations of the (approximate) simulator. Recall that the larger the value of $l$, the more accurate (and computationally expensive) such simulations will tend to be. At this point, we can re-express the objective using a telescoping sum as
@@
$\left\{\theta_{i}^{l},u_{1:m,i}^{l},\big\{G^{l}_{\theta^{l}_{i}}\big(u_{j,i}^{l}\big),G^{l-1}_{\theta^{l}_{i}}\big(u_{j,i}^{l}\big)\big\}_{j=1}^{m}\right\}\quad\text{where}\quad\theta_{i}^{l}\sim\pi,u_{1:m,i}^{l}\sim\mathbb{U},$
@@
This corresponds to an (unbiased) MLMC estimator of our original objective in ˜1. The first term, $h_{0}(\phi)$, approximates $\ell(\phi)$, but is biased since it uses $f^{0}_{\phi}$ (i.e. the lowest fidelity simulator) rather than $f^{L}_{\phi}$ (the highest fidelity simulator). The additional terms, $h_{1}(\phi),\ldots,h_{L}(\phi)$, correct this bias by estimating the expected difference between the objectives at consecutive fidelity levels.
Within each term $h_{l}(\phi)$, the functions $f_{\phi}^{l}$ and $f_{\phi}^{l-1}$ are evaluated on the same samples $u_{1:m}^{l}$ and $\theta^{l}$, i.e., they are seed-matched. This ensures that $f^{l}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l})$ and $f^{l-1}_{\phi}(u_{1:m,i}^{l},\theta_{i}^{l})$ are coupled and highly correlated, which leads to a reduction in variance (Owen, 2013, Ch. 8) since
@@
// Projecting the gradients (only when conflicting)
@@ 4 Theory
We say that $\mu$ is log-concave if it has a density of the form $\exp(-\psi(z))$ for some convex function $\psi:\mathcal{Z}\rightarrow\mathbb{R}$. We recall that for $r\in[1,\infty)$, $d,d^{\prime}\in\mathbb{N}$ and a non-empty, open, connected set $\mathcal{Z}\subseteq\mathbb{R}^{d}$, the space of vector-valued $r$-integrable functions with respect to a probability distribution $\mu$ is given by $L^{r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{L^{r}(\mu)}:=(\int_{\mathcal{Z}}\|g(x)\|_{2}^{r}\mu(dx))^{\nicefrac{{1}}{{r}}}<\infty\}$. For $\tau\in\mathbb{N}$, the corresponding Sobolev space of vector-valued functions of smoothness $\tau$ is given by $W^{\tau,r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{W^{\tau,r}(\mu)}=(\sum_{|\alpha|\leq\tau}\|D^{\alpha}g\|_{L^{r}(\mu)}^{r})^{\nicefrac{{1}}{{r}}}<\infty\}$, where for a multi-index $\alpha\in\mathbb{N}^{d}$, $D^{\alpha}$ is the weak derivative operator corresponding to $\alpha$. Finally, we recall that a function $g$ is locally $K_{\text{Lip}}$-smooth if its gradient is locally Lipschitz continuous with Lipschitz constant $K_{\text{Lip}}>0$; i.e. for all $z,z^{\prime}$ in some open set of $\mathcal{Z}$, we have that $\|\nabla g(z)-\nabla g(z^{\prime})\|_{2}\leq K_{\text{Lip}}\|z-z^{\prime}\|_{2}$. For simplicity, we will write $\tilde{q}_{\phi}(x_{1:m},\theta)$ for the conditional density model used for either NLE (in which case $\tilde{q}_{\phi}(x_{1:m},\theta)=q^{\text{NLE}}_{\phi}(x_{1}|\theta)$) and NPE (in which case $\tilde{q}_{\phi}(x_{1:m},\theta)=q^{\text{NPE}}_{\phi}(\theta|x_{1:m})$).
@@
    The generators satisfy $\left\|G^{l}\right\|_{W^{1,4}(\pi\times\mathbb{U})}\leq S_{l}\;$ for $l\in\{0,1,\ldots,L\}$.
@@
    $\log\tilde{q}_{\phi}$ is continuously differentiable, locally $K_{\text{Lip}}(\phi)$-smooth and satisfies the growth condition $\|\nabla\log\tilde{q}_{\phi}(x_{1:m},\theta)\|_{2}\leq K_{\text{grow}}(\phi)(\sum_{i=1}^{m}\|x_{i}\|_{2}+\|\theta\|_{2}+1)$ for some $K_{\text{Lip}}(\phi),K_{\text{grow}}(\phi)>0$.
@@
$\text{Var}\left[h_{0}(\phi)\right]$ $\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U})}^{4}+1\right),$
$\text{and}\;\text{Var}\big[h_{l}(\phi)\big]$ $\leq\frac{K_{l}(\phi)}{n_{l}}\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U})}.$
See Section˜A.1 for the proof, and we emphasise that the variance is over both parameters $\theta$ and noise, but conditional on $\phi$. (A1) requires log-concavity, which is a strong condition, but this is only required of the prior and the base measure. However, in SBI these tend to be simple distributions such as Gaussians or uniforms, which satisfy this assumption; see Saumard and Wellner (2014) for how to verify log-concavity in practice. (A2) is relatively mild: it asks that $G^{0},\ldots,G^{L}$ have at least one derivative in $\theta$ and $u$, and for these generators and their derivatives to have a fourth moment. This will hold when the simulators are used to define sufficiently well-behaved data-generating processes, but will be violated for sufficiently heavy-tailed distributions (e.g. the Cauchy). Finally, (A3) is mild; it holds when the gradient of the log conditional density estimator is Lipschitz continuous, which is for example the case when the conditional density is twice continuously differentiable and the Hessian is bounded (see e.g. Lemma 2.3 of Wright and Recht (2022)). It also holds for models such as mixtures of Gaussians, which are used in mixture density networks and for normalising flows with sufficiently regular transformations; see Table 1 in Liang et al. (2022) for some known Lipschitz continuous transformations. There are two key implications of Theorem˜1. The first is a bound on the variance of the MC objective:
$\text{Var}\left[{\ell}_{\text{MC}}(\phi)\right]\leq\frac{K(\phi)}{n}\left(\left\|G\right\|_{W^{1,4}(\pi\times\mathbb{U})}^{4}+1\right),$ (4)
@@
$\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U})}^{4}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U})}.$ (5)
In order to make this bound small, we need to make each term small. Typically, $\|G^{0}\|_{W^{1,4}(\pi\times\mathbb{U})}$ will be large, but this will be counterbalanced by taking $n_{0}$ to be large. For the higher-fidelity levels, the number of samples $n_{l}$ will typically be smaller, but this will be counter-balanced by the fact that $\|G^{l}-G^{l-1}\|_{W^{1,4}(\pi\times\mathbb{U})}$ is small whenever $G^{l-1}$ is a good approximation of $G^{l}$. As we now show, we can directly use ˜5 to get an indication of how to select $n_{0},\ldots,n_{L}$.